Masulis Inc. is considering a project that has the following cash flow and WACC data. What is the project's discounted payback? 

WACC: Year Cash flows 10.00% 0 1 $925 $525 2 4 3 $445 $485 $405 

  • a. 2.25 years 

  • b. 1.86 years 

  • c. 2.14 years 

  • d. 2.40 years 

  • e. 1.95 years
Answer #1

D9 f =2+46.9/334.34 A B с D E F 1 2 year 3 st Cash flows 0 $ 1 $ 2 $ 3 $ 4 $ pv@10% (925.00) 525.00 485.00 445.00 405.00 Pres

D9 fic =2+46.9/334.34 с D E F А B 1 2 year Cash flows 30 |-925 4 1 525 5 2 485 6 3 445 74 405 pv@10% 1 =C3/1.1 =C4/1.1 =C5/1.